OpenStudy (anonymous):

Could someone explain step by step, how to solve: x^2 + 3x = x^2 + 11x + 28

OpenStudy (campbell_st):

1st collect like terms subtract x^2 from both sides leaves 3x = 11x + 28

OpenStudy (anonymous):

I've gotten to that stage, but I need to find what X is, as it's a simultaneous equation.

OpenStudy (campbell_st):

its not simultaneous just subtract 3x from both sides 8x + 28 = 0 can you solve that..?

OpenStudy (anonymous):

Ah right I see, thanks! x=-7/2.

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

Yes, it helped me a lot, thanks again :)